### v3

# The taggedPBC: Annotating a massive parallel corpus for crosslinguistic investigations

Abstract
Existing datasets available for crosslinguistic investigations have tended to focus on large amounts of data for a small group of languages or a small amount of data for a large number of languages. This means that claims based on these datasets are limited in what they reveal about universal properties of the human language faculty. While this has begun to change through the efforts of projects seeking to develop tagged corpora for a large number of languages, such efforts are still constrained by limits on resources. The current paper reports on a large tagged parallel dataset which has been developed to partially address this issue. The taggedPBC contains POS-tagged parallel text data from more than 1,940 languages, representing 155 language families and 78 isolates, dwarfing previously available resources. The accuracy of particular tags in this dataset is shown to correlate well with existing SOTA taggers for high-resource languages (SpaCy, Trankit). Additionally, a novel measure derived from this dataset, the N1 ratio, correlates with expert determinations of intransitive word order in three typological databases (WALS, Grambank, Autotyp) such that a Gaussian Naive Bayes classifier trained on this feature can accurately identify basic intransitive word order for languages not in those databases. While much work is still needed to expand and develop this dataset, the taggedPBC is an important step to enable corpus-based crosslinguistic investigations, and is made available for research and collaboration via GitHub.
keywords:
datasets, POS-tagging, low resource languages, computational typology
1 Introduction
There are over 7,000 languages in the world, and such languages are rather diverse in terms of their typological features. At the same time, despite the overall diversity there are many similarities between languages, suggesting the possibility of universal properties that guide language evolution. These diversities and similarities have been identified and categorized in various ways by a variety of databases, most notably WALS [Dryer:2013ab], Grambank [Skirgard:2023ab], Autotyp [Bickel:2023aa], Glottolog [Hammarstrom:2024aa], and Lexibank [List:2022aa]. [^1] See http://wals.info ; http://grambank.clld.org ; http://github.com/autotyp/autotyp-data ; http://www.glottolog.org ; http://lexibank.clld.org There are also collections of corpora that are expansive in their coverage of individual or groups of languages yet focused on particular genres or domains (CHILDES, Brown, etc; and see the extensive corpora available via the CLARIN project). [^2] See http://childes.talkbank.org ; http://www.clarin.eu/resource-families The former tend to provide information about languages, while the latter tend to provide actual language data in various forms. The issue with the former is often lack of coverage for particular features for all languages in the dataset, whereas the latter typically suffer from insufficient annotation and lack linguistic coverage. Both kinds of datasets have a particular kind of value, but do not go far enough to address the concerns of linguistic diversity (i.e. [Benitez-Burraco:2025aa]) and gradience (i.e. [Levshina:2023ab]) needed to expand our understanding of (cross-)linguistic processes.
More recent attempts to create annotated databases that do address these issues by developing robust crosslinguistic corpora include the Universal Dependencies Treebanks project (UDT; [Zeman:2024ab]), [^3] See http://universaldependencies.org which contains data from individual languages that has been annotated for parts of speech and dependencies using a common framework. These kinds of datasets are promising in allowing for the automatic identification of properties that can categorize or classify languages in various types, which reflects an increasing interest in computational typology [Jager:2025aa], whereby annotated corpora can allow for extraction of features that align with those found in typological databases. [^4] It should be noted here that so far the focus in computational typology has been less on automatically extracting features from corpora and more on validating theories of language/universals using expert classifications found in various descriptive typological databases, a consideration that I return to in the final discussion section below. Unfortunately, such datasets still suffer many of the same issues regarding coverage and degree of annotation as identified above.
The amount of crosslinguistic coverage and degree of annotation of existing datasets are concerns when trying to investigate questions related to universal properties of human language. The reason for this is that such investigations require annotated data with information on word types or dependencies, data unavailable for most of the world’s languages. This means that crosslinguistic investigations using smaller datasets may suffer from lack of generalizability. For example, while [Hahn:2022aa] investigated dependencies in 80 languages (the largest sample of languages to date for this kind of investigation), [^5] Recent cross-linguistic investigations using tagged corpora include [Jing:2021aa] (71 languages) and [Hahn:2020aa] (51 languages). the majority of languages in their dataset were from the Indo-European language family. This calls into question the generalizability of their findings, as such findings might be biased toward a particular group of languages [Blasi:2022aa]. [^6] Similar to how other scientific findings have been biased toward WEIRD demographics [Henrich:2010aa].
In its most recent form, the UDT has 160+ hand-annotated languages available for crosslinguistic investigation, but there are some concerns with the data. For one thing, the datasets are not parallel texts, which means there is a variety of genres representing each language, such that one could question how representative they are of a particular language as well as how easily they can be compared with corpora from other languages. Additionally, the number of sentences available for each language is variable, with many languages being represented by fewer than 200 sentences in the v2.14 release. Other datasets of tagged corpora have similar issues in terms of coverage. For example, the ODIN project (and its re-packaged form via LECS) [^7] See http://depts.washington.edu/uwcl/odin ; http://huggingface.co/datasets/lecslab/glosslm-corpus contains interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences.
These datasets and related publications show the value of annotated data for crosslinguistic investigation, but they also highlight some of the difficulties inherent in creating such data. For one thing, annotated data requires a large investment in time and manpower, whereby experts in specific languages must invest time to do actual annotation of the data. Additionally, annotators and researchers must decide which features should be annotated and to what degree. This is also partly dictated by the specifics of a language (which features are relatively more rare/interesting and/or well-described) as well as the amount, quality, and kind of descriptions available for a given language. Generally the data is sourced with a particular goal or research question in mind, which then conditions the manner and kind of annotation that is done.
Recent advances in computational tools suggest the possibility of using such tools for automatic annotation, greatly reducing the time and effort needed to develop multilingual tagged datasets. As an example, there is some indication that automated taggers for under-resourced languages can benefit from high-resource language part of speech tags in parallel corpora [Agic:2015aa, Imani:2022aa], particularly for major word classes like nouns and verbs, even in zero-shot situations on low-resource languages. But while the use of such tools may greatly reduce the annotation burden for researchers, the derived dataset must also be validated for crosslinguistic investigation. Specifically, it is incumbent on researchers to identify a set of criteria used to evaluate whether the automatic annotations are sufficiently aligned with human-curated annotations or classifications before such a dataset can be considered robust enough for further crosslinguistic investigation.
The current paper reports on the development of a large corpus of parallel texts tagged for word class that partly addresses the issues identified above. With data sourced from the Parallel Bible Corpus (PBC; [Mayer:2014aa]: 1,597 languages) and additional data sourced from newly-available Bible translations (345 languages), computational word alignment tools allow for automatic tagging of word classes in 1,942 languages (roughly 25% of the world’s languages) that represent 155 families and 78 isolates.888‘Isolates’ is defined here as languages found in the corpus that do not share family membership with another language in the corpus. The automated part of speech tagging method outlined in this paper was validated by comparing the results with output from existing automatic taggers for high-resource languages (SpaCy, Trankit). [^9] See http://spacy.io/ ; http://nlp.uoregon.edu/trankit The results indicate good alignment, providing a baseline from which to conduct crosslinguistic investigations. Further, features extracted from this dataset align with word order classifications in typological databases.
Statistical properties of the derived corpus also comport with expectations regarding normality of language data. Importantly, this is partly because each language has a large amount of data available for comparison. Individual corpora in the majority of these languages (over 1,850) contain more than 1,800 parallel sentences/verses, and the smallest corpus contains more than 700 sentences/verses. This dwarfs both the linguistic data and the language diversity available from other sources. Further, the parallel nature of the data means that crosslinguistic investigations using this dataset can be confident that they are comparing similar contexts. The dataset (the taggedPBC) as well as explanations regarding its development and Python code for replicating the results described here, can be found at the linked repository. [^10] See http://github.com/lingdoc/taggedPBC
In the sections below I first outline the methodology used to develop the auto-taggers for each language (§2). I then highlight statistics of the dataset (§3) and describe the test used to assess tag validity (§3.1). I also show that this dataset allows for extraction of features associated with typological properties, identifying a novel measure derived from this dataset (the “N1 ratio”) that correlates with intransitive word order classifications in multiple typological databases (§3.2). I conclude the paper with discussion of ways the dataset could be improved to support more robust crosslinguistic investigations (§4.1) as well as potential avenues of research made possible by this dataset (§4.2).
2 Methodology
Transferring POS tags from high-resource languages has been shown to work relatively well for word classes with many members, such as Nouns and Verbs [Agic:2015aa, Imani:2022aa, Imani:2022ab]. Specifically, [Agic:2015aa] show that using an IBM Model 2 to transfer tags between aligned words in pairs of languages achieves much better results than unsupervised and weakly-supervised approaches. Additionally, [Imani:2022aa], using the PBC, were able to improve on accurate POS-tagging for low-resource languages by propagating labels from 36 high-resource languages to 17 low-resource target languages. They note that
“A limitation of the GLP [our method] is that training over a MAG (multilingual alignment graph) created for all PBC languages requires a prohibitively large amount of resources, and based on our experiments, if we use a larger number of target languages at the same time, the performance will likely drop.” (p. 1585)
While [Imani:2022aa]’s approach does result in higher overall accuracy than that of [Agic:2015aa] for specific language pairs, it is not clear that it performs better for particular word classes or for all low-resource languages more generally. Additionally, the computational resources required by Imani et al’s approach (using transformer architectures and word embeddings) are much more significant than those required by [Agic:2015aa]’s approach. Accordingly, the methodolgy used here follows that of [Agic:2015aa] in training IBM Model 2 to align words between language pairs, in this case each source language and English. It should be noted here that IBM Model 2 is a statistical word alignment model that is agnostic to the order of words, which makes it ideal for a situation like this one, where we are dealing with languages for which we don’t know the word order. The essential idea is that for each language pair we train a model to align words in a parallel corpus with two languages, then we “translate” the words in the source (low-resource) language to the target (high-resource) language, and transfer the POS-tag of the target language word to the source language word.
Another important thing to note is that while machine translation methods have expanded in recent years, there are various issues with some of the more recent (“AI”-related) approaches, particularly the use of large language models. Training or fine-tuning such models to translate between languages is not only computationally resource-intensive, but often also result in inaccuracies in output, making it questionable to transfer POS-tags. Further, developments in word alignment algorithms such as eflomal [Ostling:2016aa] and Embedding-Enhanced GIZA++ [Marchisio:2022aa] require language-specific processing (embeddings in some cases) and can also be computationally resource-intensive. Additionally, alignment algorithms such as eflomal do not handle language pairs with word order differences particularly well. In contrast, the IBM Model 2 has relatively minor computational resource requirements and is simple to implement, while depending primarily on statistical patterns that handle differences in word order between language pairs.
2.1 Improving tag accuracy
There are several additional steps that I take to improve tagging accuracy, which fall under two general principles. The first principle is related to ensuring that parallel verses share as many common concepts as possible, i.e. that they are in some sense more semantically accessible. Increasing the number of shared concepts between languages in a dataset relative to rare concepts should allow a statistical word alignment model to more successfully identify alignments between shared concepts and/or words. This is akin to “under-sampling” in machine learning, whereby relevant properties of a dataset are adjusted to reduce noise in training. The second principle is related to ensuring that specific properties of individual languages (with regards to word formation) are taken into account. Since languages differ in terms of fusional properties (how many morphemes or units of meaning a word contains), this ought to be considered by any attempt at word alignment.
Accordingly, to address the concerns present under the first principle, I reduce the number of verses in the corpus using the following criteria. I first lemmatize two English New Testament translations, then reduce the dataset to verses which share at least 4 lemmas in those translations. This ensures that verses are more likely to be semantically accessible. Translations chosen for this procedure are the New Living Translation (NLT) and the New International Version (NIV), after discussions with Bible translators who work with low-resource languages. Translators indicated that these two versions are often consulted for their readability and accessibility in terms of how they translate the original Greek text, particularly in backtranslation (a checking step before final translation verification). The smaller dataset is then further reduced to verses where a verb occurs that is also present in at least five other verses, which increases the likelihood that each individual non-English Bible translation has common concepts that can be aligned with words in the English text. This procedure results in 1,884 verses for training word alignment models. As noted above, this is a systematic method for under-sampling the parallel data, which should help improve the quality of the trained statistical word alignment model.
In order to address the second principle, I train subword tokenizers for each language. For this I leverage the MarianMT subword tokenizer, which has been developed for machine translation between multiple languages, as a base architecture. Specifically, I used the SentencePieceBPETokenizer without lowercasing (from the Huggingface Transformers library) and retrained tokenizers on each language. Training subword tokenizers for a single language allows the tokenizer to identify sub-word units that hypothetically correspond to morphemes. Using language-specific tokenizers before mapping alignments between sentences should enhance the ability of the statistical word alignment model to identify correspondences between words in the pairs of languages.
It should also be noted that the languages in the PBC use several different writing systems and scripts. Some, like Devanagari, are syllabic, while others, like Arabic, are largely phonetic, and yet others, like Chinese, are systems that use individual characters to represent words or syllables. In order to ensure that we can adequately compare languages in terms of their sound structure, and to support the training of subword tokenizers, I convert all disparate systems to romanized script using the uroman [^11] See http://github.com/isi-nlp/uroman tool, with the additional step of performing word-level tokenization for translations using scripts where words are not clearly separated by whitespace (fugashi [^12] See http://github.com/polm/fugashi for Japanese, khmernltk [^13] See http://github.com/VietHoang1512/khmer-nltk for Khmer/Kuy, jieba [^14] See http://github.com/fxsjy/jieba for Chinese, pyidaungsu [^15] See http://github.com/kaunghtetsan275/pyidaungsu for Burmese/Karen/Mon, attacut [^16] See http://github.com/PyThaiNLP/attacut for Thai, botok [^17] See http://github.com/OpenPecha/Botok for Tibetan) before converting to romanized script. [^18] As noted by a reviewer, this is not an ideal mapping, since romanization does not necessarily reflect the phonemic values of a language in the same way that a standardized orthography does. Putting aside the question of phonemic or phonetic accuracy, however, I would argue that romanization does at least allow for comparison in a way that retaining original scripts does not.
As a final step, I train each model with the source being the non-English language and the target being English. This enables a direct mapping of the words in the source language verse to the English words, allowing us to extract POS tags easily from words in the POS-tagged English sentence (via the SpaCy library). That is, once the respective translation model has been trained for a particular language, every word of every sentence/verse is back-translated to English. If the translation of that word matches a word in the respective POS-tagged English sentence/verse, the source language word is tagged with the English POS. [^19] For the purposes of cross-linguistic comparison, I employ the Universal Dependencies tag set for parts of speech which includes 6 open classes (ADJ ‘adjective’, ADV ‘adverb’, INTJ ‘interjection’, NOUN ‘noun’, PROPN ‘proper noun’, VERB ‘verb’), 8 closed classes (ADP ‘adposition’, AUX ‘auxiliary’, CCONJ ‘coordinating conjunction’, DET ‘determiner’, NUM ‘numeral’, PART ‘particle’, PRON ‘pronoun’, SCONJ ‘subordinating conjunction’) and 3 ‘other’ classes for handling characters (PUNCT ‘punctuation’, SYM ‘symbol’, X ‘other’). These 17 classes, while not used by all languages, provide enough of a basic tag set to allow for cross-linguistic comparison. This methodology enables taggging of 1,884 verses for a large number of languages in the PBC. Due to some differences in how sentences/clauses are translated between languages, and some portions not being available for some languages, there are a few mismatches. We can examine the actual number of sentences and sort them into bins, which gives counts of verses/languages in the corpus (Table 1).
Number of verses Number of languages
1800+ 1853
1500-1800 27
1000-1500 50
700-1000 12
Total 1,942
Table 1: Number of verses in each tagged language corpus
2.2 Caveats
While automated transfer of POS tags from high-resource languages such as English to low-resource languages has been shown to give decent results, as noted by the literature above, there are some necessary caveats. First, limiting the number of sentences does remove some signal along with noise, meaning that it is possible not all word types will be represented in a given language. Along with this, mapping infrequent word types or function words (i.e. that in English, der in German), which are often less prominent and more functionally diverse crosslinguistically, will likely be less succcessful. Another case in point is the presence of such terms as numeral classifiers: many languages require special words or markers that classify nouns being enumerated, but such a class of words is simply not present in other languages, like English. Since we are transferring tags from English, word classes like numeral classifiers will not be tagged in this dataset at all, which means there is scope for improving on these annotations. At the same time, the large number of parallel sentences for the vast majority of languages in this dataset, and the common method by which they were developed, means that findings for other word types (specifically nouns, proper nouns, pronouns, verbs, and auxiliaries) are statistically robust.
The methodology followed here should not be construed as a claim that all languages share the same tag set, but rather that certain word classes (nouns, verbs, etc.) exist in the majority of languages and can be readily identified via translation. At the same time, there are a large number of words in each of these corpora that simply have no tag (more than 50% of tokens in most corpora are tagged as “unk”) because there is no direct translation equivalent corresponding to a tagged English word in the same verse. This reflects the methodological goal of maximizing accuracy of POS transfer. While this leaves potentially ambiguous words for future investigation (and see the discussion in the concluding section below regarding development of language-specific automated tools or manual annotation for individual low-resource corpora), the translation method provides a decent proxy measure that supports the identification of major word classes. As I demonstrate below, despite the caveats presented here, information and features extracted from this tagged dataset comport with expected observations and lead to novel findings. This dataset is therefore released in order to provide a baseline for future research, with potential for expansion and refinement, particularly in relation to language-specific POS tags.
3 Results
The final taggedPBC dataset contains tagged data for 1,942 modern languages (1,944 if you include the 2 conlangs with NT translations: Esperanto and Klingon). This represents roughly 25% of the world’s languages from 155 distinct families and 78 single languages or isolates. The number of languages with more than 1,800 sentences/verses in the dataset is 1,853 (see Table 1), and the fewest number of verses is represented by Nafusi (ISO 639-3: jbn) with 729 verses. The automatic tagging method outlined above results in a statistically robust number of unique arguments and predicates [^20] In the current paper, I use the term “argument” and “predicate” as a way of grouping classes of words that loosely correspond to heads of phrases and serve similar roles/functions crosslinguistically. Specifically, “argument” groups the word classes of nouns, proper nouns, and pronouns, which are deemed to represent prototypical agents/patients. The term “predicate” groups the word classes of verbs and auxiliaries, which are deemed to represent prototypical actions/states. By using these terms to group POS tags, I am not claiming that these tags exclusively represent all instantiations of “arguments” or “predicates” crosslinguistically, but simply separating out those tags from other tags in the Universal Tag Set used in the overall dataset, in order to develop procedures to compute relevant statistics for crosslinguistic comparison. for all 1,942 languages (mean: 699 arguments, 466 predicates), with the fewest number of unique arguments being tagged for Nugunu (ISO 639-3: yas; 76) and the fewest number of unique predicates being tagged for Rapanui (ISO 639-3: rap; 57). If we compare the number of unique nouns and verbs, the fewest unique nouns are identified for ut-Hun (ISO 639-3: uth; 59) and the fewest unique verbs for Bambalang (ISO 639-3: bmo; 54).
As noted above, this is significantly more than currently available crosslinguistic tagged data in terms of language coverage, with the closest similar corpus-based dataset being the UDT, in which only 106 of the 160 languages in the v2.14 release have more than 200 tagged sentences. However, before we can use this dataset for investigation, we need to be confident in the validity of the tags as well as the features that can be extracted from this dataset. Accordingly, we now turn to validation of the tagging accuracy of the IBM2 word alignment POS-tag transfer method (§3.1), as well as how well extracted features correlate with existing classifications in typological databases (§3.2).
3.1 Validating tag accuracy
One way to validate the automated tags provided by the IBM2 word alignment method involves assessing against other automatic taggers developed for high resource languages. Here we want to see that the method used for tagging these languages aligns with existing language-specific taggers at a level above chance.
POS-taggers have been developed for high-resource languages, and the current state of the art is represented by the SpaCy and Trankit text processing libraries. These libraries have supervised POS-taggers for 16 and 31 languages respectively (excluding languages with non-roman scripts). For this comparison, each of the trained taggers in these libraries was used to tag the reduced PBC verses for the respective language and then the specific tags for arguments (NOUN, PROPN, PRON) and predicates (VERB, AUX) were compared with the same verses tagged via the word alignment method. [^21] I should note here, first, that both libraries use a variation of the universal tag set adopted by the UDT project and my methodology, as outlined in footnote 19. While there are some differences, they apply to less common word classes and do not affect arguments and predicates, which were the focus of this comparison. Second, these taggers were not developed using Biblical texts, which reduces the overall likelihood of correspondence, and makes the actual result more striking. Specifically, each word-alignment tagged verse was stripped of tags and re-tagged by the respective libraries. Then the two sets of tagged verses (between the library and the word-alignment method) were compared and the number of matching arguments/predicates was divided by the total number of arguments/predicates identified by the respective library in a given sentence. This was then averaged across all verses to provide an agreement percentage between the translation method and the respective library for the argument and predicate tags. [^22] See code at the linked repository for a more detailed breakdown of correspondence between individual POS tags in the argument and predicate classes. Correspondence with argument and predicate tags between the word alignment method and these pos-taggers was on average above 75% for both SpaCy (Table 2; arguments: 0.84; predicates: 0.76) and Trankit (Table 3; arguments: 0.84; predicates: 0.77).
ISO 639-3 Language Arguments Predicates
cat Catalan 0.85 0.76
hrv Croatian 0.76 0.56
dan Danish 0.87 0.84
nld Dutch 0.92 0.87
fin Finnish 0.9 0.8
fra French 0.85 0.67
deu German 0.87 0.84
ita Italian 0.87 0.8
lit Lithuanian 0.82 0.63
nob Norwegian Bokmål 0.91 0.86
pol Polish 0.82 0.65
por Portuguese 0.84 0.78
ron Romanian 0.75 0.66
slv Slovenian 0.6 0.72
spa Spanish 0.88 0.88
swe Swedish 0.9 0.79
Average 0.84 0.76
Table 2: SpaCy correspondences with the taggedPBC
ISO 639-3 Language Arguments Predicates
afr Afrikaans 0.94 0.89
eus Basque 0.69 0.78
bul Bulgarian 0.81 0.7
cat Catalan 0.84 0.88
hrv Croatian 0.76 0.7
ces Czech 0.77 0.71
dan Danish 0.86 0.9
nld Dutch 0.92 0.89
fin Finnish 0.92 0.84
fra French 0.87 0.82
deu German 0.9 0.86
hun Hungarian 0.91 0.68
ind Indonesian 0.88 0.65
gle Irish 0.82 0.46
ita Italian 0.88 0.83
lat Latin 0.86 0.73
lav Latvian 0.91 0.82
lit Lithuanian 0.88 0.76
nob Norwegian (Bokmål) 0.91 0.89
nno Norwegian (Nynorsk) 0.9 0.84
pol Polish 0.83 0.65
por Portuguese 0.86 0.8
ron Romanian 0.76 0.8
gla Scottish Gaelic 0.84 0.56
srp Serbian 0.75 0.73
slk Slovak 0.76 0.73
slv Slovenian 0.6 0.75
spa Spanish 0.89 0.9
swe Swedish 0.89 0.9
tur Turkish 0.85 0.6
vie Vietnamese 0.86 0.69
Average: 0.84 0.77
Table 3: Trankit correspondences with the taggedPBC
This comparison informs us that the translation method agrees with existing SOTA POS-taggers more than 75% of the time, which suggests that this methodology is relatively robust at transferring tags from English to other languages. At least for tagging major word classes, it aligns pretty well with automatic taggers individually developed for specific languages. However, “the proof is in the pudding”: does the data in the taggedPBC lend itself to actual correlations with observed properties of language? To assess this, we now turn to observations of word order that can be extracted from this dataset.
3.2 Testing features extracted from the dataset: word order
By “word order” most researchers refer to the relative order of referential arguments (i.e. elements referring to prototypical agents/patients, often represented by nouns, proper nouns, and pronouns) versus predicates (i.e. elements referring to prototypical actions and states as represented by verbs and auxiliaries). This makes it possible to talk about “verb-initial” vs “verb-final” or “verb-medial” languages, as well as grammatical constructs like Subject and Object in relation to the position of the Verb. In the current investigation I focus on intransitive word order, which is concerned with a predicate that requires a single core argument, resulting in three possible word orders: “verb final” (SV), “verb initial” (VS), and “free” (no dominant order). In order to identify these relations in corpora I use the following procedure for each language in the taggedPBC corpus:
  1. 1.
    Filter the verses in the corpus, retaining only verses/sentences with both an argument and a predicate.
  2. 2.
    In the resulting set, verses/sentences with an argument occurring before a predicate are classed as “noun first”, while those with a predicate occurring before an argument are classed as “verb first”. [^23] Note that the terms “noun first” and “verb first” are simply class/grouping labels (they could be termed “X” and “Y”) used to compute the N1 ratio. The terms do not constitute claims about a particular sentence structure.
  3. 3.
    This gives a count of “noun first” and “verb first” verses/sentences for each language.
  4. 4.
    Turn this into the “N1 ratio” by dividing the total number of “noun first” verses/sentences by the total number of “verb first” verses/sentences.
The resulting scalar value (the “N1 ratio”) allows us to make a plausible quantitative inference based on observation of actual text data, rather than making a categorical judgement on a language based on our knowledge of it (which may be limited or expansive depending on the particular language).
Figure 1: The N1 ratio in relation to word order in 3 typological databases
To test the validity of this measure I compared the derived values with known expert judgements of intransitive word order (SV, VS, or “free”) in three typological databases (Figure 1): the World Atlas of Language Structures (WALS; [Dryer:2013ab]), Grambank [Skirgard:2023ab], and the AUTOTYP Database [Bickel:2023aa]. For each of these databases, I conducted a one-way ANOVA with the N1 ratio as the dependent variable, and the database category as the grouping variable for all languages shared between the PBC and the respective database (WALS: 646 languages; Grambank: 937 languages; AUTOTYP: 203 languages). [^24] Code to reproduce these results can be found at the linked repository. For each comparison, the N1 ratio significantly differentiated between SV and VS languages (WALS mean diff. 1.443, p < 0.001; Grambank mean diff. 1.285, p < 0.001; AUTOTYP mean diff. 1.544, p < 0.001), and for WALS and Grambank it additionally differentiated between SV/VS and free word orders, [^25] For AUTOTYP the N1 ratio significantly differentiated between SV and free word orders (p < 0.001) but was marginal for VS and free languages (p = 0.054). indicating that this ratio can be used as a proxy for word order, at least for the PBC.
Figure 2: Distribution of the N1 ratio in the taggedPBC
As an additional verification step of the N1 ratio as a proxy for word order, I trained a machine learning classifier to predict word order classifications for languages in the PBC for which word order was unknown. Since the taggedPBC data is normally distributed (see Figure 2 and additional statistical tests at the linked repository), I trained a Gaussian Naive Bayes (GNB) classifier on the 1,152 languages with known word order (with the N1 ratio as the feature), and then predicted on the remaining 790 languages with unknown word order. This resulted in 655 of the languages being classified as “SV”, and 135 being classified as “VS”. Initial tests with other classifiers gave similar results, though some (like the Decision Tree Classifier) identified up to 10 previously unclassified languages as “free”. When literature searches were conducted for these languages (see Table 4), existing research identified all of the languages as having “VS” basic word order, validating the GNB classification.
ISO 639-3 Language GNB DTC Actual order
ifa Amganad Ifugao VS free VS
prf Paranan VS free VS
bpr Koronadal Blaan VS free VS
cnl Lalana Chinantec VS free VS
ayp North Mesopotamian Arabic VS free VS
zty Yatee Zapotec VS free VS
zar Rincón Zapotec VS free VS
mih Chayuco Mixtec VS free VS
spy Sabaot VS free VS
zca Coatecas Altas Zapotec VS free VS
Table 4: Classifications of word order by two different classifiers
This test of features indicates that the taggedPBC is a valid source of crosslinguistic comparative data. Despite potential inaccuracies in tags, features extracted directly from these corpora correlate well with expert tools and classifications, both in terms of tags in individual languages and in terms of features such as word order. This suggests some exciting new lines of research.
4 Conclusions and future work
The preceding sections have presented a new, massive dataset of parallel text tagged for parts of speech in 1,942 languages via cross-lingual POS-tag transfer, derived partially from the PBC [Mayer:2014aa] with additional data sourced from newly-accessible Bible translations. The dataset is available for research on Github as the taggedPBC, allowing for expansion/collaboration, and the tags for nouns and verbs have been validated via comparison with taggers for high-resource languages. Additionally, I have shown that the N1 ratio, a feature extracted from this dataset on the basis of arguments and predicates, differentiates between the intransitive word order (SV, VS, free) of languages, allowing for the classification of previously unclassified languages, and suggesting exciting new directions for research in computational typology.
4.1 Potential for expansion
The availability of the dataset via Github means that there are many possibilities for expansion and refinement of the data. The current release provides a baseline of POS tags via crosslingual transfer using the IBM2 model, but other alignment and transfer methods could be applied to increase accuracy for individual languages or groups of languages. As noted, there are concerns with the applicability of universal POS tags for specific languages and language families. To address this, manually tagging portions of the individual corpora could allow for the training of individual models on a per-language basis that would improve accuracy. Some guidance along these lines is presented at the linked repository to facilitate collaboration. In each case, it would be important to validate such approaches, while balancing accuracy with computational efficiency.
Individual datasets could also be expanded, either through tagging of additional portions of the Bible present in the original PBC, or by adding data from other (hand-tagged) corpora. While this would affect the overall distribution of data, it might also benefit crosslinguistic investigations in various ways. Multiple levels of annotation could also be supported, either through automatic annotation via lexical databases or manual annotation, provided that these are also validated/verified in a principled manner.
4.2 Potential research directions
This dataset provides an important source of data for the growing field of computational typology. As mentioned in footnote 4 above, this field has primarily been focused on statistical methods whereby different features of language (or of environment) are correlated with determinations or classifications in typological databases (see [Jager:2025aa]). This kind of investigation is limited by the availability of information on such features or classifications for particular languages, and there may be some debate regarding how to precisely categorize a language (with respect to word order, see the large number of “UNK” [unknown] classifications for word order in the Grambank database).
The dataset presented here, along with the example of extracting the N1 ratio and observing its relation to word order, provides a potential new direction for computational typology, whereby features can be directly extracted from corpora. As with the example above, datasets like the taggedPBC allow researchers to identify corpus-based features that correlate with (and/or differentiate between) expert classifications in typological databases. These features can then be used to classify previously unclassified languages, greatly expanding the coverage of such databases, and directly addressing the issue of linguistic diversity, as highlighted by [Benitez-Burraco:2025aa] and others. Further, corpus-derived measures allow for differing approaches to linguistic questions, such as the gradient-based word order approach proposed by [Levshina:2023ab]. While the example above shows that languages do not overlap very much in terms of the N1 ratio with regard to basic intransitive word order classifications, the point remains that no language is entirely SV or VS, supporting alternative methods to answer related questions.
In general, if we consider POS tags to be one means of identifying syntactic patterns, as suggested by the present investigation, there should be a principled, crosslinguistic way to do so. Grouping POS tags and/or observing their relative orders in sentences is one way to compute measures from corpora. It logically follows, then, that if such measures reliably correlate with other linguistic features (as the N1 ratio does with intransitive word order in WALS/Grambank/Autotyp), we should be able to use them as proxy measures for those features. Having a tagged crosslinguistic dataset of this size is valuable for answering multiple questions (such as whether a language has prepositions or postpositions), and improving the scope and quality of the annotations simply increases the kinds of questions that can be asked and answered.
Some questions that this dataset is uniquely able to support investigations in are those related to word types crosslinguistically, whether with regard to length, order, or other patterns. The fact that such questions can be queried in relation to corpora additionally allows for known effects of frequency on language [Behrens:2016aa, Bentz:2017aa, Bybee:2006aa, Haspelmath:2021aa, Koplenig:2022aa] to be accounted for. Other questions that could be asked, related to dependencies and phonological properties, may require additional annotation, whether automated, semi-automated, or manual. While improvements can definitely be made to the overall annotations, the taggedPBC dataset provides a solid baseline to push computational typology forward as a means of investigating (universal) properties of human language.